# StreamingTOM: Streaming Token Compression for Efficient Video Understanding

### v1

@@ 4.3 Efficiency Analysis
Table 3 provides a comprehensive performance profile of StreamingTOM across different configurations to guide practical deployment. Memory exhibits sublinear growth consistently due to CTR’s fixed token budget, remaining nearly constant from 16.0 GB to 16.7 GB across 16-512 frames, demonstrating bounded memory under pre-LLM compression. TTFT remains consistently low and throughput stabilizes at real-time levels across varying video lengths and batch sizes, confirming scalability for practical streaming deployment. Figure 5 breaks down the complete processing pipeline, comparing baseline and StreamingTOM across vision and query stages. These results confirm that StreamingTOM maintains efficient and predictable performance across diverse streaming scenarios.
@@ 4.4 Ablation Study
Table 4 analyzes the trade-off between per-frame detail and temporal coverage. Our default configuration (50 tokens, 4-bit) achieves the best balance: fewer tokens (40) sacrifice critical details, while more tokens (60) reduce temporal coverage under fixed memory budgets, both degrading accuracy. This configuration reduces memory footprint to 6.4% relative to the 196-token 16-bit baseline, achieving a $15.7\times$ compression ratio while maintaining strong accuracy.
@@ 5 Conclusion
This paper presents StreamingTOM, the first training-free framework that performs pre-LLM token reduction designed for streaming video understanding. Unlike existing methods that only manage post-LLM kv-cache, StreamingTOM addresses pre-LLM computation and post-LLM memory bottlenecks through a unified two-stage design. Causal Temporal Reduction (CTR) enforces causality with a fixed per-frame budget $G$, drastically reducing prefill complexity by processing only $G$ selected tokens instead of all $N$ visual tokens per frame, while Online Quantized Memory (OQM) bounds post-LLM memory through 4-bit quantization and retrieval. Experiments demonstrate that StreamingTOM achieves state-of-the-art accuracy among training-free methods, reaching 63.8% on offline benchmarks and 55.8%/3.7 on RVS with $15.7\times$ kv-cache compression. Critically, it sustains bounded memory growth over long horizons by reducing one-hour stream kv-cache from 18.8GB to 1.2GB, enabling practical deployment of video LLMs for real-time, long-duration applications.

### v2

@@ 4.3 Efficiency Analysis
Table 3 provides a comprehensive performance profile of StreamingTOM across different configurations to guide practical deployment. Memory exhibits sublinear growth consistently due to CTR’s fixed token budget, remaining nearly constant from 16.0 GB to 16.7 GB across 16-512 frames, demonstrating bounded memory under pre-LLM compression. TTFT remains consistently low and throughput stabilizes at approximately 20 tokens/s for long sequences across varying video lengths and batch sizes, confirming scalability for practical streaming deployment. Figure 5 breaks down the complete processing pipeline, comparing baseline and StreamingTOM across vision and query stages. These results confirm that StreamingTOM maintains efficient and predictable performance across diverse streaming scenarios.
@@ 4.4 Ablation Study
Table 4 analyzes the trade-off between per-frame detail and temporal coverage. Our default configuration (50 tokens, 4-bit) achieves the best balance: fewer tokens (40) sacrifice critical details, while more tokens (60) reduce temporal coverage under fixed memory budgets, both degrading accuracy. This configuration reduces memory footprint to 6.4% relative to the 196-token 16-bit baseline, achieving a $15.7\times$ compression ratio while maintaining strong accuracy. While 2-bit quantization further halves memory, it consistently yields lower accuracy across all token counts, confirming that 4-bit strikes the optimal balance between compression and fidelity.
@@ 5 Conclusion
This paper presents StreamingTOM, the first training-free framework to perform pre-LLM token reduction for streaming video understanding. Unlike existing methods that only manage post-LLM kv-cache, StreamingTOM addresses pre-LLM computation and post-LLM memory bottlenecks through a unified two-stage design. CTR enforces strict causality with a fixed per-frame budget, reducing prefill complexity, while OQM bounds post-LLM memory through 4-bit quantization and retrieval. Experiments demonstrate that StreamingTOM achieves state-of-the-art accuracy among training-free methods, reaching 63.8% on offline benchmarks and 55.8%/3.7 on RVS with $15.7\times$ kv-cache compression ratio. Critically, it sustains bounded memory growth over long horizons by reducing one-hour stream kv-cache from 18.8GB to 1.2GB, enabling deployment of video LLMs for real-time applications.
Acknowledgments
This paper is supported by Young Scientists Fund of the National Natural Science Foundation of China (NSFC) (No. 62506305), Zhejiang Leading Innovative and Entrepreneur Team Introduction Program (No. 2024R01007), Key Research and Development Program of Zhejiang Province (No. 2025C01026), Scientific Research Project of Westlake University (No. WU2025WF003), Chinese Association for Artificial Intelligence (CAAI) & Ant Group Research Fund - AGI Track (No. 2025CAAI-ANT-13). It is also supported by the research funds of the National Talent Program and Hangzhou Municipal Talent Program.
Supplementary Material
Appendix A Implementation Details
A.1 Terminology Clarification
Throughout this paper, Pre-LLM and Post-LLM refer to pipeline location, not inference stage. Pre-LLM denotes compression applied after the vision encoder and projector but before visual tokens enter the LLM transformer layers. Post-LLM denotes memory management applied after kv-cache formation within the LLM. Both operations occur during the prefill phase; the distinction captures where in the pipeline compression acts, not when during inference.
A.2 Streaming Attention and Token Saliency
CTR necessitates a saliency score for each visual token derived from the vision encoder. Rather than materializing full attention maps, we implement a streaming self-attention kernel inspired by memory-efficient chunked attention [9], integrated into the SigLIP encoder [68] to compute both the attention output and per-token importance scores simultaneously. Given queries $Q$, keys $K$, values $V$, and an optional mask $M$, the kernel processes keys in small chunks through two passes over the sequence. The first pass accumulates a numerically stable log-sum-exp of the attention scores along the key dimension, while the second pass recomputes the scores to obtain normalized probabilities. These probabilities are subsequently aggregated into the final output and an importance accumulator. Algorithm 1 summarizes this procedure.
Algorithm 1 Streaming attention with importance scores
0:  Query $Q$, keys $K$, values $V$, optional mask $M$, scale $\alpha$, chunk size $C$
0:  Attention output $O$, importance scores $s$
1:  Convert $Q,K,V$ to float32 and initialize attn_max, attn_lse, $O$, imp
2:  for each key chunk $[s:e)$ of size at most $C$ do
3:   Compute scores $S=\alpha QK_{s:e}^{\top}$
4:   if $M$ is provided then
5:    Add corresponding mask slice: $S\leftarrow S+M_{:,:,:,s:e}$
6:   end if
7:   Update attn_max and attn_lse using log-sum-exp over the key dimension
8:  end for
9:  for each key chunk $[s:e)$ do
10:   Recompute $S=\alpha QK_{s:e}^{\top}$
11:   if $M$ is provided then
12:    Add corresponding mask slice: $S\leftarrow S+M_{:,:,:,s:e}$
13:   end if
14:   Compute probabilities $P$ from attn_max and attn_lse
15:   $O\leftarrow O+PV_{s:e}$
16:   $\text{imp}_{s:e}\leftarrow\text{imp}_{s:e}+\sum_{h,q}P$
17:  end for
18:  $s\leftarrow\text{imp}/(H\cdot L_{q})$
19:  return $O,s$
Here $H$ denotes the number of attention heads and $L_{q}$ the number of query positions in the self-attention module. We enable this streaming kernel exclusively on the final layer of the vision encoder to expose the resulting importance scores as an additional one-dimensional tensor over visual tokens. CTR consumes these scores as attention-based saliency signals to rank, select, and compress visual tokens in a strictly streaming fashion without revisiting past frames.
A.3 Text–Video Retrieval
Layer-wise vision memory.
During video encoding, we extract key tensors $K^{(l)}\in\mathbb{R}^{H_{k}\times T^{(l)}\times d}$ from every transformer block prior to applying rotary embeddings to ensure they contain exclusively visual tokens. For each contiguous group of $G$ visual tokens, we compute a summary key
$k_{t}^{(l)}=\frac{1}{G}\sum_{i\in\mathcal{G}_{t}}K_{i}^{(l)},\hskip 28.80008ptt=1,\dots,\frac{T^{(l)}}{G},$
and store the resulting sequence $\{k_{t}^{(l)}\}$ alongside the exact KV cache for that layer. This process yields a layer-aligned memory bank that preserves the original attention-head dimensionality while maintaining the chronological order of visual segments.
Question-guided single-pass retrieval.
Upon receiving a question, we exclude vision placeholders and prompts to process the clean text through the language model in a single pass. At transformer layer $l$, we map the resulting query tensor $Q^{(l)}\in\mathbb{R}^{H_{q}\times T_{q}\times d}$ to the key–value head space by averaging query heads within each GQA group. Subsequently, we average across the token dimension to obtain a layer-specific question embedding
$q^{(l)}=\frac{1}{T_{q}}\sum_{j=1}^{T_{q}}Q^{(l)}_{:,j,:}\in\mathbb{R}^{H_{k}\times d}.$
Before retrieval, we flatten both $q^{(l)}$ and each $k_{t}^{(l)}$ into vectors, $\ell_{2}$-normalize them, and perform a cosine-similarity top-$K$ search within that layer. We determine the number of retrieved segments via
$K=\min\!\Bigl(\Bigl\lceil\frac{B}{G}\Bigr\rceil,\;\frac{T^{(l)}}{G}\Bigr),$
where $B$ represents the global token budget and $T^{(l)}$ denotes the total number of visual tokens stored at layer $l$. We maintain the selected indices in sorted order to ensure the temporal coherence of the retrieved segments.
Cache reconstruction for decoding.
Following this single forward pass, each layer identifies the necessary visual groups. We retrieve the corresponding KV slices and system prompt tokens from the memory bank to assemble a DynamicCache, reusing these per-layer indices throughout the decoding process. Consequently, the system incurs the question–video matching cost only once per question while permitting each layer to independently select the video segments contributing to its attention computation.
A.4 RVS Streaming Evaluation
While most benchmarks employ a conventional offline evaluation via the lmms-eval framework [70] where one video–question pair is processed at a time with full clip access, RVS operates under a strict streaming protocol. We sort all questions for a given video by their end timestamps. With frames sampled at a fixed rate such as $0.5$ fps, the model processes the video linearly and observes the content only once. Given a question ending at time $t$, we encode only the new frames arriving between the last encoding step and time $t$ to update the streaming KV cache without re-encoding earlier frames. Should a subsequent question fall within the timeframe of the previously encoded frames, we answer it using the existing cache and pass an empty image batch to indicate no additional visual input. This protocol ensures that the model generates predictions solely from video content observed up to the question time, highlighting the streaming nature of our method on RVS.
Appendix B Additional Results
B.1 LongVideoBench
Table 5: Results on LongVideoBench.
Method Setting Accuracy
LLaVA-OV-7B 32 frames 56.4
+StreamingTOM (ours) 0.5 FPS 56.3
We evaluate StreamingTOM on LongVideoBench [57], which spans a wide duration range (8s–60min). StreamingTOM retains 99.8% of the baseline accuracy (56.3 vs. 56.4). The marginal gap is expected, as LongVideoBench’s duration distribution skews shorter than VideoMME-Long and MLVU, where StreamingTOM’s streaming temporal coverage provides the greatest advantage (Table 1).
B.2 Generalization to Other VLMs
To validate backbone-agnostic generalization, we apply StreamingTOM to Qwen2.5-VL-7B-Instruct [3] on VideoMME. The core CTR and OQM mechanisms require no modification; only minor adjustments to the input format and 3D rotary position encoding are needed.
Table 6: Generalization results on Qwen2.5-VL-7B-Instruct.
Method Setting VideoMME Retain
Short Medium Long Avg
Qwen2.5-VL-7B 32 frames 72.4 61.3 50.3 61.3 100%
+StreamingTOM 32 frames 69.3 57.2 50.1 58.9 96.1%
+StreamingTOM 0.5 FPS 69.9 65.2 54.0 63.0 102.8%
At 32 frames, StreamingTOM retains 96.1% of baseline accuracy with substantially fewer tokens. Under streaming settings (0.5 FPS), temporal coverage boosts retention to 102.8%, consistent with our LLaVA-OV-7B findings (Table 1), confirming backbone-agnostic generalization without modifying the core pipeline.
Appendix C Discussion
C.1 Adaptivity of StreamingTOM
StreamingTOM capitalizes on the observation that long video sequences possess strong temporal locality and redundancy where nearby frames share content and few tokens drive downstream reasoning. CTR and OQM harness this structure within a strictly training-free paradigm. Since both modules operate exclusively on token representations and KV caches without assuming specific architectural details beyond a patch-based vision encoder and a transformer-style language model, StreamingTOM seamlessly integrates with diverse video LLMs. This compatibility applies to both offline and streaming settings and requires no modification to the original training procedure. The configurable per-frame token budget and global KV-cache budget serve as precise control mechanisms to balance accuracy, latency, and memory usage across varied deployment scenarios.
C.2 Deployment Scenarios
StreamingTOM’s bounded memory growth and predictable latency make it applicable to several practical streaming scenarios. (1) Long-duration monitoring. In applications such as surveillance and meeting assistants, the kv-cache remains bounded regardless of stream length (e.g., 1.2 GB for one hour of video), enabling continuous operation on a single GPU without memory overflow. (2) Latency-sensitive interaction. For embodied AI and autonomous driving assistants, sub-0.4s TTFT (Table 3 in the main paper) supports real-time querying over live video streams. (3) Resource-constrained deployment. The fixed per-frame budget $G$ ensures predictable compute and memory consumption per frame, facilitating capacity planning on edge devices with limited resources. Being training-free and backbone-agnostic, StreamingTOM can be integrated into existing VLM pipelines without retraining, lowering the barrier for practical adoption across diverse deployment environments.
Appendix D Limitations and Future Work
Despite the compression and efficiency gains achieved in a training-free setting, the underlying multimodal backbone fundamentally constrains the overall performance of StreamingTOM. Compression or memory management algorithms alone cannot resolve intrinsic limitations regarding fine-grained perception or long-range reasoning. Adopting stronger or task-adapted backbones would directly raise the performance ceiling, making the joint tuning of the backbone with CTR and OQM a logical extension. Furthermore, while our design optimizes prefill and bounds the active KV cache, the vision encoder remains the dominant latency source. Consequently, dense feature extraction in the visual front-end limits end-to-end acceleration even when the language model executes with high efficiency.
Several avenues exist to extend StreamingTOM and address these limitations. One path moves beyond purely training-free compression to explore lightweight tuning of CTR and OQM alongside the backbone by learning content-aware token budgets or memory allocation policies atop a frozen or partially fine-tuned video LLM. Simultaneously, reducing the computational cost of the vision encoder warrants investigation. This could involve introducing lightweight pre-encoder modules for early downsampling and motion-aware frame selection or integrating encoder-side sparsity and quantization techniques. Integrating such backbone improvements with StreamingTOM offers the potential to further minimize the end-to-end cost of long-form video understanding while preserving robust performance.